If the points (2,0),(0,1),(4,5)and (0,c)are concyclic, then the value of "c" is ?

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Shabuddin,

first find the equation of circle passes through point(2,0) ,(0,1) and (4,5)

equation of the circle

x2 +y2 +2gx +2fy +c =0

so put the points

 4+ 4g +c =0 .............1

 1 +2f +c =0 ..............2

39 + 8g + 10f +c =0 ............3

solve these equation

g =-25/12

f =-8/3

c =13/3

so equation of circle

x2 + y2 -25x/6 -16y/3 +13/3 =0

now point(0,c) satisfy this equation

c2 -16c/3 + 13/3 =0

3c2 -16c +13 =0

c=1 ,13/3

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert